Circular No. 8705

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006bq–2006bx Eight additional apparent supernovae have been found on unfiltered

CCD images: 2006bq, 2006br, and 2006bv by T. Puckett (cf. IAUC 8699; with A. Pelloni for 2006bq, G. Sostero for 2006br, and with A. Sehgal and R. Gagliano for 2006bv); 2006bs by P. Luckas, O. Trondal, and M. Schwartz (cf. IAUC 8687); 2006bt, 2006bw, and 2005bx by W. Li (cf. IAUC 8693; 2006bt with E. Lee, 2006bw with M. Baek, and 2006bx with D. R. Madison); and 2006bu by R. Holmes, H. Devore, and B. Graves (cf. IAUC 8697).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006bq Apr. 23.34 18^{h}39^{m}58.^{s}88 +39^{o}58′56′′.8 15.8 2′′.9 E, 2′′.0 N
2006br Apr. 25.23 13 30 01.80 +13 24 56.8 17.8 6′′.7 W, 1′′.1 S
2006bs Apr. 19.51 10 26 53.61 − 3 42 51.0 17.7 10′′.5 E, 3′′.9 S
2006bt Apr. 26.48 15 56 30.53 +20 02 45.6 17.5 44′′.4 W, 22′′.9 S
2006bu Apr. 27.28 13 52 47.81 + 5 18 49.4 18.1 21′′.8 E, 1′′.4 S
2006bv Apr. 28.36 12 41 01.55 +63 31 11.6 17.8 28′′.4 E, 0′′.6 S
2006bw Apr. 27.38 14 33 56.73 + 3 47 55.8 16.6 18′′.8 W, 5′′.7 N
2006bx May 1.26 10 05 14.16 +21 28 32.0 18.1 8′′.5 E, 70′′.8 N

Additional magnitudes for 2006bq in NGC 6685: 2005 Aug. 8, [19.7 (Puckett); 2006 Apr. 24.461 UT, 15.7 (A. Sehgal, Osoyoos, BC). SN 2006bq is a type-Ia supernova (just past maximum on Apr. 25.49; cf. CBET 481). Additional magnitudes for 2006br in NGC 5185: Apr. 15, [19.8 (Puckett); 25.99, 17.5 (G. Sostero, Remanzacco, Italy). SN 2006br is a type-Ia supernova (∼ 1 week before maximum on Apr. 27.36; cf. CBET 487). Additional magnitudes for 2006bs in MCG +00-27-14: Apr. 3.51, [18.0; 26.48, 17.5. SN 2006bs is a type-II supernova (∼ 2–3 weeks past maximum on Apr. 27.18; cf. CBETs 484, 487). Additional magnitudes for 2006bt: Feb. 24.51, [19.0; Apr. 27.36, 17.4. SN 2006bt is a type-Ia supernova (near maximum on Apr. 27; cf. CBET 485). Additional magnitudes for 2006bu: 1950 Apr. 19, [21.0 (Palomar Sky Survey, red); 1995 Jan. 31, [22.0 (PSS, red); 2006 Feb. 26.33, [20.3 (Holmes); Apr. 28.29, 18.1. Additional magnitudes for 2006bv in UGC 7848: Apr. 5, [19.5; 29.08, 17.8. SN 2006bv is a type-IIn supernova (near maximum on Apr. 30.14; cf. CBET 494). Additional magnitudes for 2006bw: Mar. 23.41, [18.0; May 1.40, 16.6. SN 2006bw is a type-Ia supernova (∼ 6 days past maximum on May 2.31; cf. CBET 499). Additional magnitudes for 2006bx in UGC 5434: Feb. 23.38, [19.5; May 2.22, 18.1. SN 2006bx is a type-II supernova (∼ 2–3 weeks past explosion on May 3.18; cf. CBET 500).

2006 May 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green